How is cos pi/4 + cos pi/6 different than cos (pi/4 + pi/6)? Wouldn't the second one distribute giving you the same answer?
Answer by fcabanski(1391)

In the case of trig functions what is inside the parentheses is the angle value the trig function acts upon. cos 30 degrees + cos 60 degrees is not equal to cos (30+60) = cos(90) degrees.


cos pi/4 = .707 and cos pi/6 = .866. Their sum is 1.573.


cos (pi/4 + pi/6) is cos (10pi/24) = .259.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)